OpenStudy (el_arrow):

find the area of the surface formed by revolving about the polar axis. r = 3cos(Theta) 0

OpenStudy (el_arrow):

hello

OpenStudy (el_arrow):

hey man thanks for coming to help me some this problem :)

OpenStudy (el_arrow):

*solve

OpenStudy (el_arrow):

i am still here btw

OpenStudy (anonymous):

Set \(x=r\cos\theta=3\cos^2\theta\) and \(y=r\sin\theta=3\sin\theta\cos\theta\). Then \[\begin{cases} \dfrac{dx}{d\theta}=-3\sin2\theta\\\\ \dfrac{dy}{d\theta}=3\cos2\theta \end{cases}\] The arc length of a curve \(C\) in polar form is given by \[\int_CdS=\int_C\sqrt{\left(\frac{dx}{d\theta}\right)^2+\left(\frac{dy}{d\theta}\right)^2}\,d\theta=\int_C\sqrt{r^2+\left(\frac{dr}{d\theta}\right)^2}\,d\theta\] The surface area is then given by \[\int_C2\pi r\cos\theta\,dS=2\pi\int_0^{\pi/2}3\cos^2\theta\sqrt{9\sin^22\theta+9\cos^22\theta}\,d\theta\]

OpenStudy (el_arrow):

okay and i can factor out the 9's and the sin and cos turn into a 1 right?

OpenStudy (anonymous):

Yes, the integral reduces quite a bit. \[18\pi\int_0^{\pi/2}\cos^2\theta\,d\theta\]

OpenStudy (el_arrow):

okay and what happens when it revolves about the polar axis do i use the same equation

OpenStudy (el_arrow):

this other problem is asking to find the area of the surface formed by revolving about the polar axis so i am confused about what it meant by polar axis

OpenStudy (anonymous):

The polar axis is the positive x-axis. That's already accounted for in the formula above with \(r\sin\theta\). Speaking of which, the integral actually should be \[\int_C2\pi r\color{red}{\sin}\theta\,dS=18\pi\int_0^{\pi/2}\sin\theta\cos\theta\,d\theta\] Sorry about that

OpenStudy (el_arrow):

okay so we still use that same formula of the surface area?

OpenStudy (el_arrow):

why do you have sin2(theta) and cos2(theta)?

OpenStudy (anonymous):

Yes. If you're wondering about the derivation, it's all a matter of converting from rectangular to polar. Given a curve \(C\) defined over an interval \([a,b]\) by a curve \(\begin{cases}x=x(t)\\y=y(t)\end{cases}\), the area of the surface generated by revolving the curve about the x- or y-axis, respectively, is given by \[A=2\pi\int_C y(x)\,dS\quad\text{or}\quad A=2\pi\int_Cx(y)\,dS\] where \(dS\) is the arc length over the interval, given by the expression \[dS=\sqrt{\left(\frac{dx}{dt}\right)^2+\left(\frac{dy}{dt}\right)^2}\,dt\] In converting to polar, we replace \(x=r\cos\theta\) and \(y=r\sin\theta\) and replace the derivatives accordingly with either \(\dfrac{dr}{d\theta}\) or \(\dfrac{dx}{d\theta}\) and \(\dfrac{dy}{d\theta}\). I'm using trig identities. Differentiating \(x=r\cos\theta\) with respect to \(\theta\) gives \[\frac{dx}{d\theta}=\frac{dr}{d\theta}\cos\theta-r\sin\theta\] where \(r=3\cos\theta\), and so \(\dfrac{dr}{d\theta}=-3\sin\theta\). Recall that \(\sin2x=2\sin x\cos x\), and so \[\frac{dx}{d\theta}=-3\sin\theta\cos\theta-3\cos\theta\sin\theta=-6\sin\theta\cos\theta=-3\sin2\theta\]

OpenStudy (el_arrow):

okay i see and when it says about the axis theta = pi/2

OpenStudy (el_arrow):

it says surface area formed by revolving r = e^6(theta) about the axis theta=pi/2 over the interval 0<theta<pi/2 do i go work it out the same way?

OpenStudy (anonymous):

That would be the y-axis. In that case, you only exchange \(x\) for \(y\), which is actually the first integral setup I had: \[18\pi\int_0^{\pi/2}\cos^2\theta\,d\theta\]

OpenStudy (anonymous):

So this axis, \(\theta=\dfrac{\pi}{2}\), does it apply to a different problem?

OpenStudy (el_arrow):

yes sir

OpenStudy (el_arrow):

so just switch the x and y do i also change the sin with the cos?

OpenStudy (anonymous):

That's what I mean: \(x\) corresponds to \(\cos\theta\) in the integral setup, and \(y\) to \(\sin\theta\).

OpenStudy (anonymous):

For this next problem, you'd have \[2\pi\int_0^{\pi/2}r\cos\theta\sqrt{\cdots}\,d\theta\] where \(r=e^{6\cos\theta}\).

OpenStudy (el_arrow):

okay